A research engineer has to produce nozzles with a maximum aperture diameter of 200.5mm and a minimum diameter of 60 mm and with 4 intermediate diameters such that they form a linear progression for a fluids test rig. What are the diameters? Show your answers to one decimal place.


1
Expert's answer
2022-04-06T10:31:50-0400

"a_1=60\\a_6=200.5\\d=\frac{a_6-a_1}{6-1}=28.1\\a_2=60+28.1=88.1\\a_3=88.1+28.1=116.2\\a_4=116.2+28.1=144.3\\a_5=144.3+28.1=172.4"